OpenStudy (anonymous):

What is the solution to the following system using Cramer’s Rule? 4x + 6y = 19 8x – 5y = –13

OpenStudy (anonymous):

4x + 6y = 19 8x – 5y = –13 D = -20 - 48 = -68 Dx = -95 + 78 = -17 Dy = -52 - 152 = -204 (-17/-68, -204/-68) (1/4, 3)

OpenStudy (sriram):

x=1/4 y=3

OpenStudy (anonymous):

Thanks for the medal

OpenStudy (sriram):

i cant understand what is Cramers method